OpenStudy (anonymous):

a wire is 1m long,0.2mm in diameter and has a resistance of 10omh . calculate the resistance of its material

OpenStudy (anonymous):

don't u mean the resistivity of the material?

OpenStudy (anonymous):

assuming it to be resistivity, use the following eq to get the final answer. \[R = \rho L/A\] where, R = resistance = 10 ohms L = length of the wire = 1m A = area of cross-section = pi*r^2 = 3.14*0.2*0.2*10^(-6) and \[\rho = \] resistivity = which is the unknown.

OpenStudy (kropot72):

Area of cross-section \[A=\pi \times .1\times .1\times 10^{-6}m ^{2}\]

OpenStudy (anonymous):

ah! yes....my bad! the radius will be 0.1mm...so A will be as @kropot72 says.